OpenStudy (anonymous):

HELP ME PLZZ I DONT UNDERSTAND.... :( Which function has a domain of all real numbers except x= pi/2 +or- npi? A. y= cot x B. y= 1/sin x C. y= cos x D. y= sec x

OpenStudy (anonymous):

cot = 1/cos(x), and sec = 1/cos(x) and you can't technically divide by 0 so...

OpenStudy (precal):

domain is all of the x values that is valid for the function, can't be cosx since its domain is all real numbers

OpenStudy (anonymous):

wait, cot = 1/tan(x)

OpenStudy (precal):

what is -npi?